### v2

@@ 2.1 Basic constructions and assumptions
For spaces $X$, $X^{*}$, $U$, and $W$, a subset $\Omega\subset X$, an inner solution map $S_{u}:X\to U$ and an adjoint solution map $S_{w}:X\to W$, the following hold:
@@
    We are given an inner algorithm that, on each iteration $k\geq 1$, given $\{x^{n}\}_{n=0}^{k}\subset\Omega$ and $u^{k}\in U$, produces $u^{k+1}\in U$ satisfying
    $\kappa_{u}d_{U}(u^{k+1},S_{u}(x^{k}))\leq d_{U}(u^{k},S_{u}(x^{k-1}))+\pi_{u}b_{X}(x^{k},x^{k-1})$
    for some $\pi_{u}>0$, $\kappa_{u}>1$, and distance expressions $d_{U}$ and $b_{X}$.
@@
    We are given an adjoint algorithm that, on each iteration $k\geq 1$, given $\{(x^{n},u^{n+1})\}_{n=0}^{k}\subset\Omega\times U$, and $w^{k}\in W$, produces $w^{k+1}\in W$ satisfying
    $\kappa_{w}d_{W}(w^{k+1},S_{w}(x^{k}))\leq d_{W}(w^{k},S_{w}(x^{k-1}))+{\mu_{u}}d_{U}(u^{k+1},S_{u}(x^{k}))+\pi_{w}b_{X}(x^{k},x^{k-1})$
    for some ${\mu_{u}},\pi_{w}>0$, $\kappa_{w}>1$, and a distance expression $d_{W}$.
@@
    We are given a differential transformation that, on each iteration $k\in\mathbb{N}$, given $u^{k+1}\in U$ and $w^{k+1}\in W$, produces $\widetilde{F^{\prime}}(x^{k})\in X^{*}$ that, for a target $F^{\prime}(x^{k})\in X^{*}$, satisfies for some $\alpha_{u},\alpha_{w}\geq 0$ and a distance expression $d_{X^{*}}$ the bound
@@
The inner and adjoint tracking conditions Items (i) and (ii) are parameter change aware contractivity conditions for the inner and adjoint algorithms: if $x^{k}=x^{k-1}$, the former reduces to a standard contractivity condition. The condition Item (iii) allows converting the construction error of $\widetilde{F^{\prime}}(x^{k})$ to the tracking errors of the inner and adjoint algorithms.
We next provide brief examples of inner and adjoint methods that satisfy the corresponding parts of Section 2.1. For the inner and adjoint tracking conditions Items (i) and (ii), detailed proofs are available in [29] for most of the examples presented below. To aid the reader, we explicitly include the proof for the inner forward-backward splitting method as an example.
Example 2.2 (Inner algorithm: forward-backward).
@@
for $f$ and $g$ convex in $u$, and differentiable in $(u,x)$; $\nabla f(\,\boldsymbol{\cdot}\,;x)$ $L$-Lipschitz, and $g(\,\boldsymbol{\cdot}\,;x)$ $\gamma$-strongly convex, both uniformly in $x$. This is an instance of (3) with
@@
If $S_{u}$ is Lipschitz and $\tau L\leq 1$ for a step length parameter $\tau>0$, then the forward-backward splitting updates
$u^{k+1}:=\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(u^{k}-\tau\nabla f(u^{k};x^{k}))$
satisfy Section 2.1Item (i) [29, Theorem 3.4]. Indeed, since $\tau g(\,\boldsymbol{\cdot}\,;x^{k})$ is $\tau\gamma$-strongly convex, its proximal map is $1/(1+\tau\gamma)$-contractive [33, Lemma 4.4], satisfying for all $u,\tilde{u}\in U$ that
@@
Likewise the gradient descent map $V(u):=u-\tau\nabla f(u;x^{k})$ is non-expansive when $\tau L\leq 1$, i.e., $\|V(u)-V(\tilde{u})\|\leq\|u-\tilde{u}\|.$ Since $\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(S_{u}(x^{k}))=S_{u}(x^{k})$ [7, Theorem 4.2 & Corollary 6.22], combining these inequalities yields
$(1+\tau\gamma)\|u^{k+1}-S_{u}(x^{k})\|\leq\|V(u^{k})-V(S_{u}(x^{k}))\|\leq\|u^{k}-S_{u}(x^{k})\|.$
Thus, by adding and substracting $S_{u}(x^{k-1})$, applying triangle inequality, and using the assumption that $S_{u}$ is Lipschitz, we conclude that Section 2.1Item (i) is satisfied with $\kappa_{u}=1+\tau\gamma$ and $\pi_{u}=L_{s}$ where $L_{s}$ denotes the Lipschitz constant of $S_{u}$.
@@
The Lipschitz assumption on $S_{u}$ is guaranteed in sufficiently smooth cases by the classical implicit function theorem applied to the equation $T(u,x)=0$; see [29, Appendix B]. Nonsmooth implicit function theorems and the Aubin or pseudo-Lipschitz property of the set-valued mapping $S_{u}$ are studied in, e.g., [9, 15] as well as [7, Theorem 28.3]. For the the specific case $f(u;x)=\bar{f}(u)$ and $g(u;x)=x\bar{g}(u)$ with a scalar $x$, we refer to [7, Theorem 28.5]. If $S_{u}$ has the Aubin property, it will be Lipschitz if we assume, e.g., strict convexity to ensure the uniqueness of solutions.
Example 2.4 (Inner algorithm: primal-dual).
@@
for $K\in\mathbb{L}(Z;Y^{*})$ linear and bounded to a Hilbert space $Y^{*}$, both $f$ and $g$ convex in the first parameter, differentiable in both parameters; and $g(\,\boldsymbol{\cdot}\,;x)$ $\gamma$-strongly convex uniformly in $x$. As an instance of (3), represent the Fenchel–Rockafellar primal-dual optimality conditions of this problem as the root $u$ of the mapping
@@
If $S_{u}(x)=T^{-1}(\,\boldsymbol{\cdot}\,;x)(0)$ is Lipschitz in $\Omega$, then the PDPS updates [4]
$z^{k+1}=\operatorname{prox}_{\tau f(\,\boldsymbol{\cdot}\,;x^{k})}(z^{k}-\tau K^{*}y^{k})\quad\text{and}\quad y^{k+1}=\operatorname{prox}_{\sigma g(\,\boldsymbol{\cdot}\,;x^{k})}(y^{k}+\sigma K(2z^{k+1}-z^{k}))$
satisfy Section 2.1Item (i) when $\tau\sigma\|K\|\leq 1$. The proof in [29, Theorem 3.6] is fundamentally similar to the forward-backward in Example 2.2, but requires working with operator-induced norms and monotone operators.
The next example covers discretised PDE-constrained optimisation.
Example 2.5 (Inner algorithm: linear system splitting).
For $U$ a Euclidean space and $X$ a normed space, let both $A_{x}\in\mathbb{L}(U;U)$ and $b_{x}\in U$ be Lipschitz in $x\in X$. Consider the inner constraint of $u=S_{u}(x)$ satisfying
@@
If $S_{u}(x)=A^{-1}_{x}b_{x}$ is Lipschitz in $\Omega$, then splitting $A_{x}=N_{x}+M_{x}$ per the Jacobi or Gauss–Seidel schemes, the updates $u^{k+1}=N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}u^{k})$ satisfy Section 2.1Item (i) subject to uniform versions of standard convergence conditions for these schemes [29, Examples 4.3 and 4.4]. These proofs in [29] are based on [16, Section 4.2], where squared versions of both Section 2.1Items (i) and 2.6 are proved.
The following two examples treat adjoint methods and the differential transformation for all of Examples 2.2, 2.4 and 2.5. As we have seen, $S_{u}$ for all the former examples arises from (3) for a corresponding $T$, so the next two examples can be combined with each of those examples for the satisfaction of Section 2.1Items (ii) and (iii).
Example 2.6 (Basic adjoint: linear system splitting).
With $U$ and $W_{*}$ Euclidean spaces, and $X$ a normed space, let $S_{u}$ and $T$ be given by (3), and define

### v3

@@ 2.1 The tracking assumption
Let the normed spaces $X$ and $X^{*}$ and abstract spaces $U$, and $W$, be equipped with the distance functions $b_{X}:X\times X\to[0,\infty]$, $d_{X^{*}}:X^{*}\times X^{*}\to[0,\infty]$, $d_{U}:U\times U\to[0,\infty]$, $d_{W}:W\times W\to[0,\infty]$. Then for a subset $\Omega\subset X$; an inner solution map $S_{u}:X\to U$ and an adjoint solution map $S_{w}:X\to W$; the following hold:
@@
    We are given an inner algorithm that, on each iteration $k\geq 1$, given $\{x^{n}\}_{n=0}^{k}\subset\Omega$ and $u^{k}\in U$, produces $u^{k+1}\in U$ satisfying for some $\pi_{u}>0$, $\kappa_{u}>1$ the inner trackign inequality
    $\kappa_{u}{d_{U}(u^{k+1},S_{u}(x^{k}))}\leq{d_{U}(u^{k},S_{u}(x^{k-1}))}+\pi_{u}{b_{X}(x^{k},x^{k-1})}.$
@@
    We are given an adjoint algorithm that, on each iteration $k\geq 1$, given $\{(x^{n},u^{n+1})\}_{n=0}^{k}\subset\Omega\times U$, and $w^{k}\in W$, produces $w^{k+1}\in W$ satisfying for some ${\mu_{u}},\pi_{w}>0$, $\kappa_{w}>1$ the adjoint tracking inequality
@@
    We are given a differential transformation that, on each iteration $k\in\mathbb{N}$, given $u^{k+1}\in U$ and $w^{k+1}\in W$, produces $\widetilde{F^{\prime}}(x^{k})\in X^{*}$ that satisfies for some $\alpha_{u},\alpha_{w}\geq 0$ the bound
@@
The inner and adjoint tracking conditions Items˜(i) and (ii) are parameter change aware contractivity conditions for the inner and adjoint algorithms: if $x^{k}=x^{k-1}$, the former reduces to a standard contractivity condition. For examples of such algorithms, recall Examples˜1.1 and 1.2. The condition Item˜(iii) allows converting the construction error of $\widetilde{F^{\prime}}(x^{k})$ to the tracking errors of the inner and adjoint algorithms. We provide more detailed examples of algorithms that satisfy these conditions in Section˜3.
Remark 2.3 (Distance functions).
The distance functions $d_{U}:U\times U\to\overline{\mathbb{R}}$, etc., will in the simplest cases be given by norms or semi-norms, but we do not require this, especially from $d_{U}$ and $d_{W}$. Often in Section˜6, even norms will be adapted to the algorithms under study: they will not be the standard norms of the spaces. Only $d_{X^{*}}$ will later on be required to observe a Young relationship. Even then, it will in some of our examples in Section˜6 arise from the conjugation of a semi-norm, and hence have infinite values outside a subspace. We write squared distances as $d_{U}^{2}(u,\tilde{u}):=d_{U}(u,\tilde{u})^{2}$.
Remark 2.4 (Initial iterates and first-step errors).
The initial iterates are $x^{0}$, $u^{0}$, and $w^{0}$. For the inner and adjoint variables, the initial iterates do not appear in the tracking inequalities, which start from $x^{0}$, $u^{1}$, and $w^{1}$: the idea is that each $u^{k+1}$, which is computed using $x^{k}$, is compared against the exact inner solution $S_{u}(x^{k})$, and likewise for the adjoint variable. The first-step errors $d_{U}(u^{1},S_{u}(x^{0}))$} and $d_{W}(w^{1},S_{w}(x^{0}))$}, distinct from the initialisation errors ${d_{U}}(u^{0},S_{u}(x^{0}))$ and ${d_{W}}(w^{0},S_{w}(x^{0}))$, will appear in our final estimates. They can be made small by solving the first step to a high precision. Contractive algorithms guarantee them to be smaller than the initialisation errors.
2.2 Smoothness of differential estimates
We now derive descent- and Lipschitz-type inequalities for the approximate differential $\widetilde{F^{\prime}}(x^{k})$}, extending these classical smoothness concepts to account for differential errors under the tracking framework. Most of these result are straightforward consequences of Section˜2.1 and the scalar recurrence results of Appendix˜A.
We recall that if $F^{\prime}$ is $L$-Lipschitz, it then satisfies the descent inequality [7, Theorem 7.1]
(8) $\langle F^{\prime}(x^{k})|x-x^{k}\rangle_{X^{*},X}\geq F(x)-F(x^{k})-\frac{L}{2}{d_{X}^{2}(x,x^{k})}\quad\text{for all}\quad x,x^{k}\in X$
for the choice ${d_{X}^{2}(x,x^{k})}=\|x-x^{k}\|_{X}$. We allow here any distance expression ${d_{X}}:X\times X\to\overline{\mathbb{R}}$ such that $d_{X^{*}}$} is Young to $d_{X}$}. By this we mean that both expressions are one-homogeneous and
In particular, we allow $d_{X}$} and $d_{X^{*}}$} to be given by (semi-)norms generated by operators. We treat the descent inequality with respect to them in Section˜5.
If ${b_{X}}={d_{X}}$, simply taking $\bar{x}=x^{k}$ and $\tilde{\gamma}=\varsigma_{p}$ in the theorem to follow, and combining with the descent inequality (8), we obtain the inexact descent inequality at $x=x^{k+1}$,
(9) $\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{L+2\varsigma_{p}}{2}{d_{X}^{2}(x^{k+1},x^{k})}-\frac{1}{2\tilde{\gamma}}e_{p,k}.$
According to the theorem, the final error term has a bounded sum, which we will be able to manage in convergence proofs of optimisation methods (in Section˜4).
In such convergence proofs, it is frequently convenient to directly use the three-point descent inequality (see [7, Corollary 7.2] for the convex case, or [32, Appendix B] for the non-convex case)
(10) $\langle F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq F(x)-F(\bar{x})+\frac{\beta}{2}{d_{X}^{2}(x,\bar{x})}-\frac{L}{2}{d_{X}^{2}(x,x^{k})}\quad\text{for all}\quad x,x^{k},\bar{x}\in X.$
If, again, ${b_{X}}={d_{X}}$, combining the next theorem with this inequality, we obtain the inexact version at $x=x^{k+1}$,
(11) $\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq F(x^{k+1})-F(\bar{x})+\frac{\beta-\tilde{\gamma}}{2}{d_{X}^{2}(x^{k+1},\bar{x})}-\frac{\varsigma_{p}^{2}\tilde{\gamma}^{-1}+L}{2}{d_{X}^{2}(x^{k+1},x^{k})}-\frac{1}{2\tilde{\gamma}}e_{p,k}.$
We write
(12) $\kappa:=\min(\kappa_{u},\kappa_{w})\quad\text{and}\quad\overline{\kappa}:=\max(\kappa_{u},\kappa_{w}).$
Theorem 2.5.
Suppose Section˜2.1 holds, $X$ is a normed space, $X^{*}$ its dual with $d_{X^{*}}$} Young to $d_{X}$}. Also let $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\geq 1$, and pick $p\in[1,\kappa)$. Then, for any $\tilde{\gamma}>0$ and $\bar{x}\in X$, we have
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq-\frac{\tilde{\gamma}}{2}{d_{X}^{2}(x^{k+1},\bar{x})}-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}{b_{X}(x^{k+1},x^{k})}-\frac{1}{2\tilde{\gamma}}e_{p,k}$
for some $\varsigma_{p}\geq 0$ and $e_{p,k}\in\mathbb{R}$ that satisfies $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}<\infty$.
Proof 2.6.
By the Young relationship of $d_{X}$} and $d_{X^{*}}$},
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq-\frac{1}{2\tilde{\gamma}}{d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}-\frac{\tilde{\gamma}}{2}{d_{X}^{2}(x^{k+1},\bar{x})}.$
Let ${\varrho_{k}}:={b_{X}(x^{k},x^{k-1})}$, ${d^{u}_{k}}:={d_{U}(u^{k},S_{u}(x^{k-1}))}$, ${d^{w}_{k}}:={d_{W}(w^{k},S_{w}(x^{k-1}))}$, and ${\widetilde{d}_{k}}={d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}$. Then Section˜2.1 and $\{x^{n}\}_{n=0}^{k}\subset\Omega$ imply Appendix˜A for the index $k$. Now an application of Lemma˜A.5 establishes
${d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\leq\varsigma_{p}^{2}{b_{X}(x^{k+1},x^{k})}+e_{p,k},$
where $\varsigma_{p}$ and $e_{p,k}$ defined in Eqs.˜67 and 72 satisfy the claimed bounds by the lemma. Combining these two inequalities establishes our claim.
Remark 2.7.
From Eqs.˜67 and 72 we can obtain the more precise bounds
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}\leq\frac{{d_{U}^{2}(u^{1},S_{u}(x^{0}))}}{\pi_{u}}\bigg(\frac{\varsigma_{p}\alpha_{u}\kappa}{\kappa-1}+\frac{\varsigma_{p}\alpha_{w}{\mu_{u}}}{(\kappa-1)^{2}}\bigg)+\frac{{d_{W}^{2}(w^{1},S_{w}(x^{0}))}}{\pi_{w}}\bigg(\frac{\varsigma_{p}\alpha_{w}\kappa}{\kappa-1}\bigg)$
and
(13) $\varsigma_{p}\leq\frac{(\alpha_{u}\pi_{u}+\alpha_{w}\pi_{w})\kappa\overline{\kappa}}{p(\kappa-p)}+\frac{\alpha_{w}{\mu_{u}}\pi_{u}\overline{\kappa}}{p^{2}(\kappa-p)^{2}}.$
The sum of the error terms $e_{p,k}$ can, therefore, be made arbitrarily small by taking good-quality first inner and adjoint steps.
It will also be useful to have a Lipschitz-type property. If $d_{X^{*}}$ satisfies the triangle inequality, combining the next theorem with $F^{\prime}$ being $L$-Lipschitz with respect to the distances $d_{X^{*}}$} and $d_{X}$}, we can get the Lipschitz property with error for $\widetilde{F^{\prime}}$,
${d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x))}\leq L{d_{X}(x,x^{k})}+\sqrt{e_{\mathrm{lip},k}}\qquad(x\in X).$
According to the theorem, the error terms again have bounded sum, if the squares of residual terms $b_{X}(x^{k+1},x^{k})$} from the tracking inequalities have a bounded sum. Many standard convergence proofs, include the ones of Section˜4 will generally ensure the latter.
Theorem 2.8.
Suppose Section˜2.1 holds, and that $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\geq 1$. Then
(14) ${d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\leq e_{\mathrm{lip},k},$
where $e_{\mathrm{lip},k}\geq 0$ is such that
(15) $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty\quad\text{whenever}\quad\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}{b_{X}^{2}(x^{k+1},x^{k})}<\infty.$
Proof 2.9.
Let ${\varrho_{k}}:={b_{X}(x^{k},x^{k-1})}$, ${d^{u}_{k}}:={d_{U}(u^{k},S_{u}(x^{k-1}))}$, ${d^{w}_{k}}:={d_{W}(w^{k},S_{w}(x^{k-1}))}$, as well as ${\widetilde{d}_{k}}={d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}$. Then Section˜2.1 and $\{x^{n}\}_{n=0}^{k}\subset\Omega$ imply Appendix˜A for the index $k$. Now Lemma˜A.3 with $p=1$ gives (14) for $e_{\mathrm{lip},k}:=\breve{e}_{1,k}$ as defined in (68). The limiting property (15) is a consequence of Lemma˜A.7.
3 Inner and adjoint algorithms
We next provide brief examples of inner and adjoint methods that satisfy the corresponding parts of Section˜2.1; hence iterative ways to construct differential estimates $\widetilde{F^{\prime}}(x^{k})$ that satisfy the inexact smoothness-type properties of Section˜2.2. This section is largely based on [30], with some streamlined proofs, and some omitted proofs. The framing of the proofs in our differential estimation framework, is obviously new.
3.1 Inner algorithms and the inner tracking inequality
Forward-backward splitting
@@ 3.1 Inner algorithms and the inner tracking inequality
for $f$ and $g$ convex in $u$, and differentiable in $(u,x)$. This is an instance of (1) with
@@
Theorem 3.1.
Suppose $\nabla f(\,\boldsymbol{\cdot}\,;x)$ $L$-Lipschitz, and $g(\,\boldsymbol{\cdot}\,;x)$ $\gamma$-strongly convex, both uniformly in $x\in\Omega$ for an $\Omega\subset X$ If $S_{u}$, as defined above, is Lipschitz in $\Omega$, and $\tau L\leq 1$ for a step length parameter $\tau>0$, then Section˜2.1Item˜(i) holds with $\kappa_{u}=1+\tau\gamma$ and $\pi_{u}=L_{S_{u}}$ for the inner forward-backward splitting updates
$u^{k+1}:=\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(u^{k}-\tau\nabla f(u^{k};x^{k})).$
Proof 3.2.
This is essentially [30, Theorem 3.4], but we replicate the proof here as an example. With $k\geq 1$, let $x^{k},x^{k-1}\in\Omega$. Since $\tau g(\,\boldsymbol{\cdot}\,;x^{k})$ is $\tau\gamma$-strongly convex, its proximal map is $1/(1+\tau\gamma)$-contractive [34, Lemma 4.4], satisfying for all $u,\tilde{u}\in U$ that
@@
Likewise the gradient descent map $V(u):=u-\tau\nabla f(u;x^{k})$ is non-expansive when $\tau L\leq 1$, i.e., $\|V(u)-V(\tilde{u})\|\leq\|u-\tilde{u}\|$. Since $\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(V(S_{u}(x^{k})))=S_{u}(x^{k})$ [7, Theorem 4.2 & Corollary 6.22], combining these inequalities yields
$\begin{split}(1+\tau\gamma)\|u^{k+1}-S_{u}(x^{k})\|&\leq\|V(u^{k})-V(S_{u}(x^{k}))\|\leq\|u^{k}-S_{u}(x^{k})\|\\ &\leq\|u^{k}-S_{u}(x^{k-1})\|+\|S_{u}(x^{k})-S_{u}(x^{k-1})\|\\ &\leq\|u^{k}-S_{u}(x^{k-1})\|+L_{S_{u}}\|x^{k}-x^{k-1}\|.\end{split}$
@@
The Lipschitz assumption on $S_{u}$ is guaranteed in sufficiently smooth cases by the classical implicit function theorem applied to the equation $T(u,x)=0$; see [30, Appendix B]. Nonsmooth implicit function theorems and the Aubin or pseudo-Lipschitz property of the set-valued mapping $S_{u}$ are studied in, e.g., [10, 17] as well as [7, Theorem 28.3]. If $S_{u}$ has the Aubin property, it will be Lipschitz if we assume, e.g., strict convexity to ensure the uniqueness of solutions. For the specific case $f(u;x)=\bar{f}(u)$ and $g(u;x)=x\bar{g}(u)$ with a scalar $x$, we refer to [7, Theorem 28.5].
Primal-dual proximal splitting
@@
for $K\in\mathbb{L}(Z;Y^{*})$ linear and bounded to a Hilbert space $Y^{*}$, both $f$ and $g$ convex in the first parameter, and differentiable in both parameters. As an instance of (1), we represent the Fenchel–Rockafellar primal-dual optimality conditions of this problem as a root $u$ of the mapping
@@
Theorem 3.4.
Suppose $g(\,\boldsymbol{\cdot}\,;x)$ is $\gamma$-strongly convex uniformly in $x\in\Omega$ for a $\Omega\subset X$. If $S_{u}(x)=T^{-1}(\,\boldsymbol{\cdot}\,;x)(0)$ is Lipschitz in $\Omega$, and the step length parameters $\tau,\sigma>0$ satisfy $\tau\sigma\|K\|\leq 1$, then Section˜2.1Item˜(i) holds for the inner PDPS updates
$z^{k+1}=\operatorname{prox}_{\tau f(\,\boldsymbol{\cdot}\,;x^{k})}(z^{k}-\tau K^{*}y^{k})\quad\text{and}\quad y^{k+1}=\operatorname{prox}_{\sigma g(\,\boldsymbol{\cdot}\,;x^{k})}(y^{k}+\sigma K(2z^{k+1}-z^{k})).$
The proof in [30, Theorem 3.6] is fundamentally similar to the forward-backward splitting in Theorem˜3.1, but requires working with operator-induced norms and monotone operators.
Linear system splitting
We now cover algorithms for PDEs as the inner problems. For $U$ a Euclidean space and $X$ a normed space, let both $A_{x}\in\mathbb{L}(U;U)$, modelling a linear PDE parametrised $x$, and the right hand side $b_{x}\in U$ be Lipschitz in $x\in X$. Consider the inner constraint of $u=S_{u}(x)$ satisfying
@@
Two solve (16) inexactly and efficiently, we split $A_{x}$ into two components, as in standard Gauss–Seidel or Jacobi splittings.
Assumption 3.4 (Admissible splitting).
Let $X$ be a normed space and $U$ a Hilbert space, and $A_{x}\in\mathbb{L}(U;U)$ for all $x$ in a set $\Omega\subset X$. We assume to be given splittings $A_{x}=N_{x}+M_{x}$ with $N_{x}$ invertible, and such that $\|N^{-1}_{x}M_{x}\|\leq\zeta$ for some $\zeta\in[0,1)$.
Theorem 3.5.
Suppose Section˜3.1 holds and that $S_{u}(x)=A^{-1}_{x}b_{x}$ is $L_{S_{u}}$-Lipschitz in $\Omega$. Then the updates $u^{k+1}=N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}u^{k})$ satisfy Section˜2.1Item˜(i) in $\Omega$ with $\kappa_{u}=\zeta^{-1}$ and $\pi_{u}=L_{S_{u}}$.
Proof 3.6.
We have
$u^{k+1}-S_{u}(x^{k})=N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}u^{k})-N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}S_{u}(x^{k}))=-N^{-1}_{x^{k}}M_{x^{k}}(u^{k}-S_{u}(x^{k})).$
Hence
$\zeta^{-1}\|u^{k+1}-S_{u}(x^{k})\|\leq\|u^{k}-S_{u}(x^{k})\|\leq\|u^{k}-S_{u}(x^{k-1})\|+L_{S_{u}}\|x^{k+1}-x^{k}\|.$
To provide examples of Section˜3.1, we extend the condition to splittings that are admissible for the adjoint equations that we treat next.
Assumption 3.6 (Adjoint admissible splitting).
Let $X$ be a normed space and $U$ a Hilbert space with $A_{x}\in\mathbb{L}(U;U)$ for all $x$ in a set $\Omega\subset X$. We assume to be given splittings $A_{x}=N_{x}+M_{x}$ with $N_{x}$ invertible, and such that
(17) $\|N^{-1}_{x}M_{x}\|\leq\zeta\quad\text{and}\quad\gamma_{N}\|N^{-1}_{x}\|\leq 1$
for some $\zeta\in[0,1)$ and $\gamma_{N}>0$.
Example 3.7 (Jacobi splitting [30, Example 4.3]).
Let $A_{x}\in\mathbb{R}^{n\times n}$, and take $N_{x}$ as the diagonal of $A_{x}$. Regarding Section˜3.1, the first part of (17) reduces to strict diagonal dominance, see [13, §10.1]. The second part always holds and $N_{x}$ is invertible when the main diagonal of $A_{x}$ has only positive entries. Then $\gamma_{N}$ is the minimum of the diagonal values.
Example 3.8 (Gauss–Seidel splitting [30, Example 4.4]).
Let $A_{x}\in\mathbb{R}^{n\times n}$, and take $N_{x}$ as the upper triangle and diagonal of $A_{x}$. Regarding Section˜3.1, the first inequality of (17) holds for some $\zeta\in[0,1)$ when $A_{x}$ is symmetric and positive definite. This follows similarly to [13, proof of Theorem 10.1.2]. The second inequality holds for some $\gamma_{N}$ when $N_{x}$ is invertible, i.e., has no zeros on the main diagonal.
3.2 Adjoint algorithms and the adjoint tracking inequality; the differential transformation condition
We now treat adjoint methods and the differential transformation, i.e., Appendix˜AItems˜(i) and ii. Here, $S_{u}$ arises from (1) for a corresponding $T$, so the next two approaches to the adjoint equation and differential transformation can be combined with each of inner algorithms of the previous subsection.
Basic adjoint: linear system splitting
With $U$, $W_{*}$, and $X$ normed spaces, let $S_{u}$ and $T$ be given by (1), and define